# Chasing Anharmonicities in Polarization-Orientation Raman Spectra of Acene Crystals with Machine Learning

### v1

@@ IV.3 Temperature Dependence of PO-Raman Signals
In panels c–f of Fig. 4, a clear redshift and broadening of all peaks is observed as the temperature increases in both crystals. The extracted peak positions are listed in Table 1. The shift is similar across different modes, spanning a range of cm-1. This effect is a direct result of the incorporation of anharmonicity, in particular the temperature-dependent lattice expansion. As shown in Section S4 of SM 111See Supplemental Material at [URL will be inserted by publisher] for details on normal modes projections, supercell convergence, perpendicular configuration PO-maps, Raman tensors benchmarking and full harmonic PO-Raman results. for anthracene, the redshift is almost negligible when lattice expansion is not taken into account, despite the anharmonic dynamics. This highlights the strong influence of lattice parameter variations on the dynamics of low-frequency intermolecular modes. Notably, some modes even exhibit a blue shift in the absence of lattice expansion. We will analyze the temperature dependence of these PO-patterns in more detail in the following sections, after scrutinizing different sources of PO-pattern modification.
@@
IV.3.1 Accounting for nuclear quantum effects
Low-frequency vibrations such as the ones we have discussed in the previous sections are typically well-modeled by classical nuclear motion. The reason is that thermal energy is larger than the energy-spacing between any two vibrational levels of these modes, even at low temperatures ( K cm-1), making classical Boltzmann statistics a good approximation in these cases. In addition, the effective mass of the modes is also large, making classical dynamics equally appropriate at the temperatures (100 to 295 K) and frequency-ranges (20 to 150 cm-1) we are concerned with.
However, mode coupling between these low-frequency intermolecular vibrations and higher frequency modes at the region of cm-1 are known to be present in these crystals [57, 58]. Because these higher-frequency modes are more sensitive to nuclear quantum effects, this coupling can lead to non-trivial temperature-dependence that is sensitive to the quantum nature of the nuclei [59, 60].
In order to investigate these effects, we trained a path-integral coarse-grained (PIGS) MACE MLIP, following Refs.[61, 41]. The elevated temperature was chosen to be 500 K. We performed “elevated-temperature” centroid molecular dynamics (-CMD) [41, 62] simulations for the PO Raman spectrum of anthracene at 100 K and 295 K using the -PIGS method [41]. The results are shown in Fig. 5. As can be seen by comparing the PO-maps in Fig. 5a) and c) with the ones presented in Fig. 4a) and e), the polarization patterns in the intermolecular vibrations region are unaltered by nuclear quantum effects. At higher frequencies, changes are observed. Comparing classical and quantum unpolarized spectra, as shown in Fig. 5b-d, a typical red-shift induced by the inclusion of zero-point motion is observed already at the lower-frequency intramolecular vibrations around 750 cm-1 (highlighted in the insets of panels b-d). Importantly, the temperature dependent changes of the low-frequency modes are not altered by the inclusion of NQE.
While cross-correlations between different vibrational modes could be affected by NQE also in the low-frequency region, these play a minor role in the analysis of PO-Raman spectra, as detailed below. Therefore, we conclude that classical-nuclei simulations are sufficient to obtain anharmonic PO-Raman signals of these molecular crystals at the low-frequency range.
IV.3.2 Changes to PO-Raman Patterns Within a Quasi-Harmonic Approximation
As a first step towards the analysis of the impact of temperature on the polarization dependence of the Raman intensity, we compute the harmonic PO-dependence (parallel configuration) of the Raman intensity in anthracene and naphthalene with Eq. 7, accounting for thermal lattice expansion at 100 K and 295 K, as shown in Fig. 6 for the and modes as an example case. This approach is commonly referred to as quasi-harmonic approximation. For completeness, we show the patterns for all modes in parallel and perpendicular configurations in the SM 111See Supplemental Material at [URL will be inserted by publisher] for details on normal modes projections, supercell convergence, perpendicular configuration PO-maps, Raman tensors benchmarking and full harmonic PO-Raman results. (Section S6). In this subsection, when we refer to different temperatures, we are simply referring to the corresponding different lattice constants.
The polarization pattern, i.e the angular dependence of the Raman intensity for a given mode in Eq. 7, is fundamentally determined by the shape, symmetry, and the relative magnitude and sign of the components of the corresponding . We provide explicit expressions for the general angular dependencies of the Raman intensity patterns generated by and tensors in Section S6 of SM 111See Supplemental Material at [URL will be inserted by publisher] for details on normal modes projections, supercell convergence, perpendicular configuration PO-maps, Raman tensors benchmarking and full harmonic PO-Raman results.. In the case of naphthalene and anthracene, with the crystals oriented in the crystallographic plane (001), we show how, once the overall intensity is normalized at each temperature, no variations with lattice expansion can occur in the polarization patterns of modes, both in the parallel and perpendicular configuration within the harmonic approximation of Eq. 7. The same is true for modes in the perpendicular configuration. Only modes in the parallel configuration can be affected by lattice expansion when the relative value of the Raman tensor components of these modes is altered. As a note, we demonstrate in the SI (Section S5 of SM 111See Supplemental Material at [URL will be inserted by publisher] for details on normal modes projections, supercell convergence, perpendicular configuration PO-maps, Raman tensors benchmarking and full harmonic PO-Raman results.), that even small inaccuracies in the normal modes can lead to Raman tensors that yield incorrect polarization dependencies. These changes may be overlooked when the variable is integrated out, making PO-patterns a much stricter test for the quality of the theoretical framework compared to simply computing unpolarized spectra.
We start by discussing the PO-patterns of the anthracene crystal. Like in the fully anharmonic case shown in Fig. 4, all modes show a four-fold oscillating pattern across polarization angles, as exemplified in Fig. 6a, and the modes show more variability (see also Fig. S6 of SM 111See Supplemental Material at [URL will be inserted by publisher] for details on normal modes projections, supercell convergence, perpendicular configuration PO-maps, Raman tensors benchmarking and full harmonic PO-Raman results.). Specifically, and have and elements of opposite sign, resulting in a four-fold pattern with two maxima of different intensities, whereas displays a two-fold oscillation due to having these diagonal elements of the same sign. This is an important observation, as despite the fact that a square modulus appears in Eq. 7, the relative sign of the Raman tensor components is qualitatively relevant in the resulting PO-dependence of modes, and can therefore be extracted from a fit of the experimental data. Variations of the PO-patterns at different temperatures are minimal. As a matter of fact, we observe a change in the relative values of the diagonal components of the mode, resulting in a change of intensity of the relative maximum at 270∘ in Fig. 6a.
In the case of naphthalene, the same analysis holds true for the modes, which show a four-fold oscillating pattern, as expected from the fact that these crystals belong to the same space group. The modes again show either a two-fold or a four-fold pattern with maxima of different intensities, depending on the sign of the diagonal components of their Raman tensors (see Fig. S7 of SM 111See Supplemental Material at [URL will be inserted by publisher] for details on normal modes projections, supercell convergence, perpendicular configuration PO-maps, Raman tensors benchmarking and full harmonic PO-Raman results.). For example, an inspection of the Raman tensor of mode of naphthalene at 100 K explains both the weak PO-dependence of this mode in the parallel configuration that we observed in Fig. 4 and the absence of PO-dependence in the perpendicular configuration (see Section S3 of SM 111See Supplemental Material at [URL will be inserted by publisher] for details on normal modes projections, supercell convergence, perpendicular configuration PO-maps, Raman tensors benchmarking and full harmonic PO-Raman results.), as the and components are almost equal in magnitude and sign.
Quite surprisingly, though, in naphthalene we observe pronounced changes in the overall polarization patterns of modes with thermal lattice expansion. This is particularly evident in the case of mode , where a significant change in the relative values of the diagonal components of the Raman tensor leads to a transition from a four-fold pattern with two maxima of different intensities at 100 K to a two-fold pattern at 295 K (see Fig. 6b). This observation highlights that even within the quasi-harmonic approximation, temperature-induced lattice expansion can lead to substantial modifications in the polarization dependence of certain Raman active modes. In Section S6 of SM 111See Supplemental Material at [URL will be inserted by publisher] for details on normal modes projections, supercell convergence, perpendicular configuration PO-maps, Raman tensors benchmarking and full harmonic PO-Raman results., we show, by comparing the normal modes displacement vectors at the two different temperatures, that the and modes of naphthalene present a measurable change in direction with the different lattice parameters, leading to significant variations in their Raman tensor components and consequently in their polarization patterns. This is not the case for anthracene, where the normal modes obtained for the structures at different temperatures overlap almost perfectly with each other.
In summary, once the experimental configuration is fixed, only the PO-dependence of modes of a certain symmetry can be affected by thermal lattice expansion in the quasi-harmonic approximation. In the case of anthracene and naphthalene, these are the modes in the parallel configuration. The extent to which the PO-pattern of these modes is affected varies significantly between different materials, as exemplified by the comparison between anthracene and naphthalene. Therefore, while we prove that not all temperature-dependent changes in PO-Raman pattern stem from dynamical coupling between different vibrational modes, we also provide a clear baseline for the analysis of temperature-dependent spectra. The changes that we can capture within the quasi-harmonic approximation, although pronounced, can still be fitted by a second-rank tensor of fixed symmetry. Any additional changes in the polarization patterns must originate from anharmonic dynamical effects involving the vibrational degrees of freedom.
IV.3.3 The Impact of Mode Mixing and Mode Broadening on Anharmonic Spectral Weights
In order to start addressing the role of anharmonicity in the vibrational dynamics of anthracene, we analyze the vibrational density of states (VDOS) and the cross-VDOS (CVDOS) computed in the normal mode basis for the -point intermolecular modes. These quantities are defined in Eq. 16 of Section VI. The VDOS provides a measure of the distribution of vibrational frequencies, while the CVDOS captures correlations between different modes, revealing signatures of mode coupling. One can think of the RGDOS method discussed in Sec. III.2 as a weighting of the -point VDOS and CVDOS with Raman tensors, thus modifying the intensity ratio of the peaks and introducing the angular dependence.
In Fig. 7 we show the VDOS and CVDOS at the point of anthracene at 100 K and 295 K. For completeness, we include the same figure for naphthalene in the SM 111See Supplemental Material at [URL will be inserted by publisher] for details on normal modes projections, supercell convergence, perpendicular configuration PO-maps, Raman tensors benchmarking and full harmonic PO-Raman results. (Section S8). At low temperatures, the VDOS of anthracene shows clear Lorentzian-like peaks for each mode. The corresponding CVDOS shows hints of mode coupling between modes -, - and -. The CVDOS of modes with different symmetry is zero and not shown.
As the temperature increases, we observe the emergence of spectral weight in the VDOS of a given mode stemming from other modes. This is particularly evident for the - case at 295 K (inset of Fig. 7b) and provides a clear indication of anharmonic mode mixing. The phenomenon of mode mixing has been addressed in the analysis of Benshalom et al. [10]. They showed that in a two-mode model, the presence of anharmonic coupling leads to significant modifications in the frequency dependence of the Raman response. Instead of isolated Lorentzian peaks, the spectral functions acquire contributions from multiple modes, resulting in nontrivial intensity redistribution and the possibility of temperature-dependent changes in the PO-patterns at a given frequency.
IV.3.4 Contrasting Simulations with Experimental Data
@@
Before focusing on the angle dependence of the integrated intensity, we highlight some general discrepancies between our simulated and the experimental spectra of anthracene. We observe good overall agreement with the experimental peak positions as reported in Fig. S7 of the SI of Ref. [8], with redshifts in the order of 5-10 cm-1 in the simulated spectra. Notably, the mode has negligible Raman intensity in the experiments, where no polarization dependence could be extracted. In contrast, the mode shows appreciable intensity in our results. This discrepancy is attributed to the level of theory employed, and is not affected by the subsequent dynamics. In fact, it is known to be already present when calculating a harmonic unpolarized Raman spectrum with the same level of theory (PBE+MBD), as reported in Ref. [8].
@@
In this work, we attempt to extract the PO-dependence of the integrated intensity of every peak from the anharmonic PO-maps of Fig. 4 by fitting the Raman spectra of anthracene at 100 K, 220 K and 295 K to a multi-Lorentzian function at each angle. The fitting procedure mimics the one employed for the experiments, as detailed in the SI of Ref. [8], and reported in Section VI. In analyzing the outcome of this procedure, we should keep in mind its shortcomings stemming from the mode broadening effects discussed in Section IV.3.3.
@@
In Fig. 8 we show the fitted PO-dependence of the integrated intensity of the low-frequency Raman peaks of anthracene in both the parallel and perpendicular experimental configurations at 100, 220 and 295 K. The fitted data is normalized to the maximum intensity of each peak at the corresponding temperature. For comparison, we also include the harmonic PO-dependence and the experimental results from Ref. [8], at the closest available temperatures of 80 K, 220 K and 290 K. We remind the reader that for anthracene, the harmonic PO-pattern is insensitive to lattice expansion. In Fig. S10 of the SM 111See Supplemental Material at [URL will be inserted by publisher] for details on normal modes projections, supercell convergence, perpendicular configuration PO-maps, Raman tensors benchmarking and full harmonic PO-Raman results., we show the result of attempting the same fit to the simulated spectra of naphthalene, which was only partially successful for the lowest temperatures, due to the massive broadening at larger temperatures.
@@
The fitted intensity oscillation patterns from the simulated data in the parallel configuration shown in Fig. 8a are in general agreement with the experimental results, with the notable exception of the mode. While both theory and experiment agree that the temperature variation has no impact on the PO-dependence of this mode, its polarization pattern differs. In experiment, it shows a four-fold pattern, with two relative maxima. In the simulations, only a two-fold pattern is present. This discrepancy is entirely determined by the form of the Raman tensor predicted by the specific level of theory employed, and cannot attributed to the subsequent dynamics. Specifically, our PBE Raman tensor has and components of the same sign, which cannot give rise to a four-fold pattern. One must note that experimental artifacts, such as birefringence, can alter the polarization dependence of the Raman intensity in experiment, making the direct comparison even more challenging. Artificially including birefringence in the form of Jones matrices[63] can alter the simulated PO-patterns of modes, but not to a sufficient extent to explain the observed discrepancies (see Section 7 of SM 111See Supplemental Material at [URL will be inserted by publisher] for details on normal modes projections, supercell convergence, perpendicular configuration PO-maps, Raman tensors benchmarking and full harmonic PO-Raman results.). In the perpendicular configuration, all modes present a four-fold pattern in our simulations. This is partially consistent with experiments, where indeed every mode appears as four-fold, but with having two relative maxima of different intensities.
The experimental results in Fig. 8 show a clear change of the PO-dependence with temperature of the and modes in the parallel configuration. The largest change is seen in going from 220 K to 290 K. A similar effect is present for the mode in experiment, but less pronounced. In all cases, the effect can be described as a gradual ”loss” of the oscillation pattern with increasing temperature. When comparing to our simulated, anharmonic PO-maps, we do not find a one-to-one, quantitative correspondence of this phenomenon. In some cases the trend is reproduced correctly, such as the mode, where the oscillation pattern is indeed less pronounced at higher temperatures in both experiment and theory in the parallel configuration. In the case of the and modes, which show a strong loss of polarization dependence in experiment, our simulated pattern is not affected by temperature. Nonetheless, we observe variations of the PO pattern in the mode, with a consistent trend of gradual attenuation of the oscillation pattern with increasing temperature. Given the analysis in Section IV.3.3, we would expect that modes and would both show the effect of temperature due to their coupling, but we only observe changes in . These observations further underline the difficulties of interpreting the fitted PO-Raman maps.
@@
In our simulations, the mode shows variations in the PO pattern with temperature, although we attribute a portion of this variation to the change in the relative magnitudes of the diagonal components of the Raman tensor, discussed in Section IV.3.2 and shown in Fig. 6. In the perpendicular configuration of Fig. 8b, and still show subtle deviations in the simulated patterns. In contrast, the experimental data is substantially less clear regarding the temperature dependence of the PO patterns when the perpendicular geometry is employed. In this experimental geometry, changes are still observed for and , but no monotonic trend can be identified with the increasing temperature.
@@
Our results suggest that, while genuine mode coupling like the one that emerges in Fig. 7 can induce real changes in the polarization dependence, the additional overlap of non-Lorentzian, broadened peaks can also lead to apparent intensity sharing between modes, complicating the interpretation of PO-pattern evolution with temperature. Modes like and , which show a pronounced change in the PO-pattern in experiment, both overlap substantially with an intense peak of a different symmetry: and , respectively. Especially at polarization angles where one of the two overlapping features is in principle not present, any deviation of the other mode spectral feature from a perfect Lorentzian will lead to intensity sharing, thus potentially hiding real anharmonic coupling effects or giving rise to artificial contributions.

### v2

@@ IV.2 Temperature Dependence of PO-Raman Signals
In panels c–f of Fig. 3, a clear redshift and broadening of all peaks is observed as the temperature increases in both crystals. The extracted peak positions are listed in Table 1. The shift is similar across different modes, spanning a range of cm-1. This effect is a direct result of the incorporation of anharmonicity, in particular the temperature-dependent lattice expansion. As shown in Section S7 of SM 111See Supplemental Material at [URL will be inserted by publisher] for details on normal modes projections, supercell convergence, perpendicular configuration PO-maps, Raman tensors benchmarking and full harmonic PO-Raman results. for anthracene, the redshift is almost negligible when lattice expansion is not taken into account, despite the anharmonic dynamics. This highlights once more the strong influence of lattice parameter variations on the dynamics of low-frequency intermolecular modes. Notably, some modes even exhibit a blue shift in the absence of lattice expansion. Beyond these classical anharmonic effects, we also assessed the role of nuclear quantum effects (NQEs) on the PO-patterns. As shown in Section S8 of SM 111See Supplemental Material at [URL will be inserted by publisher] for details on normal modes projections, supercell convergence, perpendicular configuration PO-maps, Raman tensors benchmarking and full harmonic PO-Raman results. (see additional Refs. [69, 45, 51, 68] therein), based on path-integral molecular dynamics simulations of the PO-Raman patterns of anthracene at 100 K and 295 K [79, 57, 20], NQEs are found to have a negligible impact on the polarization patterns of these modes below 400 cm-1, confirming that classical-nuclei simulations are very accurate at the frequency range considered here.
In the following subsections, we will compare our simulated spectra with available experimental data and analyze in detail the signatures of anharmonic mode mixing and peak broadening that emerge at higher temperatures.
@@
IV.2.1 Contrasting Simulations with Experimental Data
@@
Before focusing on the angle dependence of the integrated intensity, we highlight some general discrepancies between our simulated and the experimental spectra of anthracene. We observe good overall agreement with the experimental peak positions as reported in Fig. S7 of the SM of Ref. [4], with redshifts in the order of 5-10 cm-1 in the simulated spectra. Notably, the mode has negligible Raman intensity in the experiments, where no polarization dependence could be extracted. In contrast, the mode shows appreciable intensity in our results. This discrepancy is attributed to the level of theory employed, and is not affected by the subsequent dynamics. In fact, it is known to be already present when calculating a harmonic unpolarized Raman spectrum with the same level of theory (PBE+MBD), as reported in Ref. [4]. In the SM, Section S13, we further show that using different functionals to calculate the crystal vibrations shift the position of the peaks in a non-uniform fashion in this frequency range, but do not alter the relative peak intensities.
@@
In this work, we attempt to extract the PO-dependence of the integrated intensity of every peak from the anharmonic PO-maps of Fig. 3 by fitting the Raman spectra of anthracene at 100 K, 220 K and 295 K to a multi-Lorentzian function at each angle. The fitting procedure mimics the one employed for the experiments, as detailed in the SM of Ref. [4], and reported in Section VI.
@@
In Fig. 4 we show the fitted PO-dependence of the integrated intensity of the low-frequency Raman peaks of anthracene in both the parallel and perpendicular experimental configurations. The fitted data is normalized to the maximum intensity of each peak at the corresponding temperature. For comparison, we also include the harmonic PO-dependence and the experimental results from Ref. [4], at the closest available temperatures of 80 K, 220 K and 290 K. We remind the reader that for anthracene, the harmonic PO-pattern is insensitive to lattice expansion. In Fig. S13 of the SM 111See Supplemental Material at [URL will be inserted by publisher] for details on normal modes projections, supercell convergence, perpendicular configuration PO-maps, Raman tensors benchmarking and full harmonic PO-Raman results., we show the result of attempting the same fit to the simulated spectra of naphthalene, which fails at higher temperatures due to excessive broadening of the peaks.
@@
The fitted intensity oscillation patterns from the simulated data in the parallel configuration shown in Fig. 4a are in general agreement with the experimental results, with the notable exception of the mode. While both theory and experiment agree that the temperature variation has no impact on the PO-dependence of this mode, its polarization pattern differs. In experiment, it shows a four-fold pattern, with two relative maxima. In the simulations, only a two-fold pattern is present. This discrepancy is entirely determined by the form of the Raman tensor predicted by the specific level of theory employed, and cannot attributed to the subsequent dynamics. Specifically, our PBE Raman tensor has and components of the same sign, which cannot give rise to a four-fold pattern. One must note that experimental artifacts, such as birefringence, can alter the polarization dependence of the Raman intensity in experiment, making the direct comparison even more challenging. Artificially including birefringence in the form of Jones matrices [47] can alter the simulated PO-patterns of modes, but not to a sufficient extent to explain the observed discrepancies (see Section S9 of SM 111See Supplemental Material at [URL will be inserted by publisher] for details on normal modes projections, supercell convergence, perpendicular configuration PO-maps, Raman tensors benchmarking and full harmonic PO-Raman results.). In the perpendicular configuration, all modes present a four-fold pattern in our simulations. This is partially consistent with experiments, where indeed every mode appears as four-fold, but with having two relative maxima of different intensities.
@@
In our simulations, the mode shows variations in the PO pattern with temperature, although we attribute a portion of this variation to the change in the relative magnitudes of the diagonal components of the Raman tensor, discussed in Section IV.1 and shown in Fig. 2.
The experimental results in Fig. 4 show a clear change of the PO-dependence with temperature of the and modes in the parallel configuration. The largest change is seen in going from 220 K to 290 K. A similar effect is present for the mode in experiment, but less pronounced. In all cases, the effect can be described as a gradual loss of the oscillation pattern with increasing temperature. When comparing to our simulated, anharmonic PO-maps, we do not find a one-to-one, quantitative correspondence of this phenomenon. In some cases the trend is reproduced correctly, such as the mode, where the oscillation pattern is indeed less pronounced at higher temperatures in both experiment and theory in the parallel configuration. In the case of the and modes, which show a strong loss of polarization dependence in experiment, our simulated pattern is not affected by temperature. Nonetheless, we observe variations of the PO pattern in the mode, with a consistent trend of gradual attenuation of the oscillation pattern with increasing temperature. In the perpendicular configuration of Fig. 4b, still shows subtle deviations in the simulated patterns. In contrast, the experimental data is substantially less clear regarding the temperature dependence of the PO patterns when the perpendicular geometry is employed. In this geometry, changes are still observed in experiment for and , but no monotonic trend can be identified with the increasing temperature. We proceed with a deeper theoretical analysis of these observations.
IV.2.2 The Impact of Mode Mixing and Mode Broadening on Anharmonic Spectral Weights
In order to address the role of anharmonicity in the vibrational dynamics of anthracene, we analyze the vibrational density of states (VDOS) and the cross-VDOS (CVDOS) computed in the normal mode basis for the -point intermolecular modes. These quantities are defined in Eq. 16 of Section VI. The VDOS provides a measure of the distribution of vibrational frequencies, while the CVDOS captures correlations between different modes, revealing signatures of mode coupling.
In Fig. 5 we show the VDOS and CVDOS at the point of anthracene at 100 K and 295 K. For completeness, we include the same figure for naphthalene in the SM 111See Supplemental Material at [URL will be inserted by publisher] for details on normal modes projections, supercell convergence, perpendicular configuration PO-maps, Raman tensors benchmarking and full harmonic PO-Raman results. (Section S10). Panels a-b display the mode-projected VDOS, whereas panels c–f report the CVDOS between all pairs of modes with the same symmetry.
At low temperatures, the VDOS of anthracene shows clear Lorentzian-like peaks for each mode. The corresponding CVDOS shows hints of mode coupling between modes -, - and -. The CVDOS of modes with different symmetry is zero and not shown. As the temperature increases, we observe the emergence of spectral weight in the VDOS of a given mode stemming from other modes. This is particularly evident for the - case at 295 K (inset of Fig. 5b) and provides a clear indication of anharmonic mode mixing. Note how in the CVDOS of Fig. 5d, the cross-correlation intensity between these two modes vanishes going from 100 K to 295 K. Instead, the VDOS gains spectral weight for the and modes, signaling that the phonon basis is a less accurate representation of the vibrational dynamics at higher temperature. This observation can explain the attenuation of the PO-pattern of the mode observed in the simulated data of Fig. 4. The phenomenon of mode mixing has also been addressed in the analysis of Benshalom et al. [12] mentioned at the beginning of the previous section. They showed how, instead of isolated Lorentzian peaks, the spectral functions acquire contributions from multiple modes, resulting in nontrivial intensity redistribution and the possibility of temperature-dependent changes in the PO-patterns at a given frequency.
@@ IV.2.2 The Impact of Mode Mixing and Mode Broadening on Anharmonic Spectral Weights
This analysis suggests that, while genuine mode coupling like the one that emerges in Fig. 5 can induce real changes in the polarization dependence, the additional overlap of non-Lorentzian, broadened peaks can also lead to apparent intensity sharing between modes, complicating the interpretation of PO-pattern evolution with temperature. Modes like and , which show a pronounced change in the PO-pattern in experiment (Fig. 4b), both overlap substantially with an intense peak of a different symmetry: and , respectively. Especially at polarization angles where one of the two overlapping features is in principle not present, any deviation of the other mode spectral feature from a perfect Lorentzian will lead to intensity sharing, thus potentially hiding real anharmonic coupling effects or giving rise to artificial contributions.
Finally, we note how the RGDOS method explained in Sec. III.2 amounts to a weighting of the -point VDOS and CVDOS with Raman tensors. This operation adds an intensity modulation that is angle-resolved, and can highlight or suppress mode coupling at different polarization angles. Therefore, one can also perform a mode decomposition of the PO-maps of Fig. 3 in terms of self- and cross-mode contributions at each polarization angle. We present in Section S12 of SM 111See Supplemental Material at [URL will be inserted by publisher] for details on normal modes projections, supercell convergence, perpendicular configuration PO-maps, Raman tensors benchmarking and full harmonic PO-Raman results. examples of this decomposition for anthracene at selected angles, at both temperatures. While such analysis provides a powerful tool to aid the deconvolution of experimental PO maps, the conclusions drawn from the analysis of the VDOS and CVDOS remain unaltered.